OpenStudy (anonymous):

]What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

Did u try it?

jhonyy9 (jhonyy9):

x-y=-3 --- y=x+3 x+3y=5 x+3(x+3)=5 x+3x+9=5 4x=-4 x= -4/4 = -1 x=-1